# QH9: A Quantum Hamiltonian Prediction Benchmark for QM9 Molecules

### 22 Jan 2024 v3

@@ Abstract
Supervised machine learning approaches have been increasingly used in accelerating electronic structure prediction as surrogates of first-principle computational methods, such as density functional theory (DFT). While numerous quantum chemistry datasets focus on chemical properties and atomic forces, the ability to achieve accurate and efficient prediction of the Hamiltonian matrix is highly desired, as it is the most important and fundamental physical quantity that determines the quantum states of physical systems and chemical properties. In this work, we generate a new Quantum Hamiltonian dataset, named as QH9, to provide precise Hamiltonian matrices for 999 molecular dynamics trajectories and 130,831 stable molecular geometries, based on the QM9 dataset. By designing benchmark tasks with various molecules, we show that current machine learning models have the capacity to predict Hamiltonian matrices for arbitrary molecules. Both the QH9 dataset and the baseline models are provided to the community through an open-source benchmark, which can be highly valuable for developing machine learning methods and accelerating molecular and materials design for scientific and technological applications. Our benchmark is publicly available at https://github.com/divelab/AIRS/tree/main/OpenDFT/QHBench.
@@ 1 Introduction
To provide a realistic dataset and comprehensive evaluation, we generate a new quantum tensor dataset named QH9. This dataset contains Hamiltonian matrices for 130,831 stable molecular geometries and 999 molecular dynamic trajectories. In order to provide comprehensive studies for quantum tensor networks, we have designed four specific tasks. The first two tasks, QH9-stable-id and QH9-stable-ood, aim to explore the performance of the networks in both in-distribution and out-of-distribution scenarios, specifically focusing on stable molecular geometries. The QH9-dynamic-geo task follows the setting of the mixed MD17, containing the same molecule with different geometries in the training, validation, and test. On the other hand, the QH9-dynamic-mol task splits the trajectories based on different molecules. Finally, we evaluate the transferability of the trained models on molecules with larger sizes, thereby testing the models’ ability to generalize beyond the training dataset. To demonstrate the quality of the predicted Hamilton matrix, we use four metrics. These metrics are based on the Mean Absolute Error (MAE) of the predicted Hamiltonian matrix $\bm{H}$, as well as the derived properties such as orbital energies $\bm{\epsilon}$ and electronic wavefunction $\psi$. Furthermore, to evaluate the quality of the predicted Hamiltonian in accelerating DFT calculations, we calculate the DFT optimization ratio by taking the model predictions as DFT initialization. The target and content of the proposed QH9 dataset and benchmark are demonstrated in Figure 1.
@@ 3.1 Datasets
Dataset Generation. For the QH9 dataset, we use open-source software PySCF (Sun et al., 2018, 2020) to conduct computational quantum chemistry calculations. In the QH9, there are two sub datasets. The first one is the QH9-stable dataset containing Hamiltonian matrices for 130,831 molecules with their geometries. We obtain the molecules and their geometries in QH9-stable from the QM9 version with 130,831, which is widely used in molecular property prediction tasks in the literature (Schütt et al., 2018; Gasteiger et al., 2020; Liu et al., 2022; Wang et al., 2022). Note that we cover all the molecules in this QM9 version. The second one is the QH9-dynamic dataset, it has molecular trajectories for $999$ molecules and each trajectory contains $100$ geometries. To obtain the accurate Hamiltonian matrices for this dataset, we set the hyper-parameters of the DFT algorithms to a tight level. Specifically, we set the grid density level to 3 to calculate accurate electronic density, and the SCF convergence condition is set to SCF tolerance of $10^{-13}$ and gradient threshold of $3.16\times 10^{-5}$ to ensure the the final states achieve tight convergence. For the density functional, we select the B3LYP exchange-correlation functional to conduct DFT calculations, and GTO orbital basis Def2SVP is selected to approximate the electronic wavefunctions. To accelerate and achieve the convergence of SCF algorithm, we use DIIS algorithm with consideration of the 8 previous steps. For the QH9-dynamic dataset, molecular dynamics simulations are conducted under the microcanonical ensemble, where the number of particles, volume, and energy remain constant (NVE). The temperature is set to 300K, and the time step for recording the molecular trajectory is set to $0.12$ fs with $1,000$ total steps. In the dataset collection, we sampled $100$ time steps for each trajectory with a timestep of $1.2$ fs. The generated QH9 dataset is available at Zenodo (https://zenodo.org/records/8274793) and GitHub (https://github.com/divelab/AIRS/tree/main/OpenDFT/QHBench/QH9).
(a) The histogram of the molecule size for QH9-stable dataset.
(b) Number of heavy atom percentage in QH9-stable dataset.
(c) The histogram of the molecule size for QH9-dynamic.
Figure 2: The dataset statistics on QH9-stable and QH9-dynamic, including molecule size distribution and percentage of molecules with different number of heavy atoms.
Dataset Statistics. The statistical data, including the number of molecules and geometries for QH9-stable and QH9-dynamic, is presented in Table 1. These molecules consist of no more than 9 heavy atoms and are composed of four specific heavy atoms: carbon (C), nitrogen (N), oxygen (O), and fluorine (F). The distribution of molecule size for QH9-stable and QH9-dynamic is shown in Figure 1(a) and Figure 1(c). Meanwhile, the percentage of molecules in QH9-stable with different the number of heavy atoms is shown in 1(b).
@@ 3.2 Tasks
QH9-dynamic-geo $99,900$ $999$ $79,920/9,990/9,990$
QH9-dynamic-mol $99,900$ $999$ $79,900/9,900/10,100$
@@
QH9-dynamic-geo. For this split and the following QH9-dynamic-mol split, there are $999$ molecular dynamics trajectories, while each trajectory includes $100$ geometries. In QH9-dynamic-geo, the split is performed geometry-wise. Specifically, for each molecule, $100$ geometries are randomly divided into $80$ for training, $10$ for validation, and $10$ for testing. Here, the molecules in the test set are visible during training but the geometric structures are different from training structures.
QH9-dynamic-mol. In this QH9-dynamic-mol split, the $999$ molecules are divided into training, validation, and testing subsets in a ratio of $0.8/0.1/0.1$. Importantly, different from the above QH9-dynamic-geo setup, all $100$ geometries corresponding to a specific molecule are grouped together and assigned to the same subset. This setup introduces a more challenging task than QH9-dynamic-geo since the geometries in the testing set correspond to different molecules as those in training.
@@ 4 Experiments
QH9-dynamic-geo QHNet $149.62$ $92.88$ $96.85$ $834.47$ $94.45$
QH9-dynamic-mol QHNet $416.99$ $153.68$ $173.92$ $9719.58$ $79.15$
@@
Geometry-wise vs. molecule-wise generalization. We further explore geometry-wise and molecule-wise generalizability by analyzing the difficulty differences between the QH9-dynamic-geo and QH9-dynamic-mol tasks. We consider the results in Table 2 for these two tasks to be comparable given that both models are trained with a similar number of geometry structures. We note that the model in the QH9-dynamic-geo task demonstrates numerically better test performance than the model in the QH9-dynamic-mol task. This is consistent with our intention when designing the tasks. Specifically, in QH9-dynamic-geo, although the geometric structures in the test set are different, the molecules themselves are not entirely novel to the model due to the exposure during the training phase. In comparison, the QH9-dynamic-mol task presents a more challenging and demanding scenario. In particular, the test set geometries in QH9-dynamic-mol correspond to entirely different molecules than those seen during training. This task requires the model to generalize from its learned patterns to the unseen molecular structures. To summarize, both tasks serve as valuable testbeds in evaluating the model’s generalization ability, and our analysis shows that QH9-dynamic-mol task, which requires extrapolating to entirely new molecular structures, is notably more challenging and demanding.
Table 4: The performance of DFT calculation acceleration. Both models, trained on the QH9-stable-id split and the QH9-stable-ood split respectively, are evaluated on a common set of $50$ randomly chosen molecules from the intersection of their test sets, making their results directly comparable. Similarly, results from models trained on the QH9-dynamic-geo and QH9-dynamic-mol splits can be compared directly.
@@
QH9-dynamic-geo 1e Optimal ratio $0.056\scriptstyle{{{\color[rgb]{0.3,0.3,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,0.3}\pgfsys@color@rgb@stroke{0.3}{0.3}{0.3}\pgfsys@color@rgb@fill{0.3}{0.3}{0.3}\pm 0.006}}}$ QH9-dynamic-mol 1e Optimal ratio $0.056\scriptstyle{{{\color[rgb]{0.3,0.3,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,0.3}\pgfsys@color@rgb@stroke{0.3}{0.3}{0.3}\pgfsys@color@rgb@fill{0.3}{0.3}{0.3}\pm 0.006}}}$
@@ Appendix A Appendix
Statistic Type QH9-stable-id QH9-stable-ood QH9-dynamic-geo QH9-dynamic-mol
# of atoms Mean 18.02 / 18.02 / 18.03 16.97 / 21.25 / 23.67 16.53 / 16.53 / 16.53 16.56 / 16.38 / 16.38
Median 18 / 18 / 18 17 / 21 / 23 17 / 17 / 17 17 / 17 / 17
Max 29 / 29 / 29 20 / 22 / 29 19 / 19 / 19 19 / 19 / 19
Min 3 / 6 / 4 3 / 21 / 23 10 / 10 / 10 10 / 10 / 10
# of electronics Mean 65.89 / 65.90 / 65.86 64.98 / 68.83 / 70.57 64.68 / 64.68 / 64.68 64.71 / 64.44 / 64.68
@@
Max 74 / 74 / 74 74 / 74 / 74 74 / 74 / 74 74 / 70 / 70
Min 10 / 18 / 24 10 / 56 / 58 38 / 38 / 38 40 / 38 / 46
Size of Hamiltonian Mean 141.59 / 141.62 / 141.56 138.99 / 149.85 / 155.09 137.96 / 137.96 / 137.96 138.04 / 137.34 / 137.88
Median 144 / 144 / 144 142 / 150 / 154 140 / 140 / 140 142 / 142 / 142
Max 166 / 166 / 166 148 / 152 / 166 146 / 146 /146 146 / 146 / 146
Min 18 / 36 / 48 18 / 126 / 130 82 / 82 / 82 86 / 82 / 92
@@ A.1 Detailed dataset statistics
A.2 Test performance over molecule size
@@
A.3 Test performance over various training sets

### 21 Mar 2024 v4

@@ Abstract
Supervised machine learning approaches have been increasingly used in accelerating electronic structure prediction as surrogates of first-principle computational methods, such as density functional theory (DFT). While numerous quantum chemistry datasets focus on chemical properties and atomic forces, the ability to achieve accurate and efficient prediction of the Hamiltonian matrix is highly desired, as it is the most important and fundamental physical quantity that determines the quantum states of physical systems and chemical properties. In this work, we generate a new Quantum Hamiltonian dataset, named as QH9, to provide precise Hamiltonian matrices for 999 or 2998 molecular dynamics trajectories and 130,831 stable molecular geometries, based on the QM9 dataset. By designing benchmark tasks with various molecules, we show that current machine learning models have the capacity to predict Hamiltonian matrices for arbitrary molecules. Both the QH9 dataset and the baseline models are provided to the community through an open-source benchmark, which can be highly valuable for developing machine learning methods and accelerating molecular and materials design for scientific and technological applications. Our benchmark is publicly available at https://github.com/divelab/AIRS/tree/main/OpenDFT/QHBench.
@@ 1 Introduction
To provide a realistic dataset and comprehensive evaluation, we generate a new quantum tensor dataset named QH9. This dataset contains Hamiltonian matrices for 130,831 stable molecular geometries and 999 or 2998 molecular dynamic trajectories. In order to provide comprehensive studies for quantum tensor networks, we have designed four specific tasks. The first two tasks, QH9-stable-id and QH9-stable-ood, aim to explore the performance of the networks in both in-distribution and out-of-distribution scenarios, specifically focusing on stable molecular geometries. The QH9-dynamic-geo task follows the setting of the mixed MD17 (Yu et al., 2023), containing the same molecule with different geometries in the training, validation, and test. On the other hand, the QH9-dynamic-mol task splits the trajectories based on different molecules. Based on the number of geometries within the QH9-dynamic datasets, it is divided into QH9-dynamic-100k for 999 trajectories and QH9-dynamic-300k for 2998 trajectories. Finally, we evaluate the transferability of the trained models on molecules with larger sizes, thereby testing the models’ ability to generalize beyond the training dataset. To demonstrate the quality of the predicted Hamilton matrix, we use four metrics. These metrics are based on the Mean Absolute Error (MAE) of the predicted Hamiltonian matrix $\bm{H}$, as well as the derived properties such as orbital energies $\bm{\epsilon}$ and electronic wavefunction $\psi$. Furthermore, to evaluate the quality of the predicted Hamiltonian in accelerating DFT calculations, we calculate the DFT optimization ratio by taking the model predictions as DFT initialization. The target and content of the proposed QH9 dataset and benchmark are demonstrated in Figure 1.
@@ 3.1 Datasets
Dataset Generation. For the QH9 dataset, we use open-source software PySCF (Sun et al., 2018, 2020) to conduct computational quantum chemistry calculations. In the QH9, there are two sub datasets. The first one is the QH9-stable dataset containing Hamiltonian matrices for 130,831 molecules with their geometries. We obtain the molecules and their geometries in QH9-stable from the QM9 version with 130,831, which is widely used in molecular property prediction tasks in the literature (Schütt et al., 2018; Gasteiger et al., 2020; Liu et al., 2022; Wang et al., 2022). Note that we cover all the molecules in this QM9 version. The second one is the QH9-dynamic dataset, it has molecular trajectories for $999$ molecules and each trajectory contains $100$ geometries. To obtain the accurate Hamiltonian matrices for this dataset, we set the hyper-parameters of the DFT algorithms to a tight level. Specifically, we set the grid density level to 3 to calculate accurate electronic density, and the SCF convergence condition is set to SCF tolerance of $10^{-13}$ and gradient threshold of $3.16\times 10^{-5}$ to ensure the the final states achieve tight convergence. For the density functional, we select the B3LYP exchange-correlation functional to conduct DFT calculations, and GTO orbital basis Def2SVP is selected to approximate the electronic wavefunctions. To accelerate and achieve the convergence of SCF algorithm, we use DIIS algorithm with consideration of the 8 previous steps. For the QH9-dynamic dataset, molecular dynamics simulations are conducted under the microcanonical ensemble, where the number of particles, volume, and energy remain constant (NVE), and the temperature is set to 300K. For QH9-dynamic-100k, the time step for recording the molecular trajectory is set to $0.12$ fs with $1,000$ total steps and then sampled one step every 10 steps. For QH9-dynamic-300k, it conducts MD simulations with time step of $1.2$ fs with $100$ total steps. The generated QH9 dataset is available at Zenodo (https://zenodo.org/records/8274793) and GitHub (https://github.com/divelab/AIRS/tree/main/OpenDFT/QHBench/QH9).
Dataset Statistics. The statistical data, including the number of molecules and geometries for QH9-stable and QH9-dynamic, is presented in Table 1. These molecules consist of no more than 9 heavy atoms and are composed of four specific heavy atoms: carbon (C), nitrogen (N), oxygen (O), and fluorine (F). The detailed statistics are shown in Table 5.
@@ 3.2 Tasks
QH9-dynamic-100k-geo a $99,900$ $999$ $79,920/9,990/9,990$
QH9-dynamic-100k-mol $99,900$ $999$ $79,900/9,900/10,100$
QH9-dynamic-300k-geo b $299,800$ $2,998$ $239,840$/$29,980$/$29,980$
QH9-dynamic-300k-mol $299,800$ $2,998$ $239,800$/$29,900$/$30,100$
  • a
    QH9-dynamic-100k conducts MD simulations with time step of $5$ a.u. ($\approx 0.120944$ fs) for $1,000$ total steps, and selects one sample every 10 steps.
  • b
    QH9-dynamic-300k conducts MD simulations with time step of $50$ a.u. ($\approx 1.20944$ fs) for 100 total steps.
Table 1: The statistics of our defined four tasks.
@@
QH9-dynamic-geo. For this split and the following QH9-dynamic-mol split, they use the molecular dynamics trajectories, while each trajectory includes $100$ geometries. In QH9-dynamic-geo, the split is performed geometry-wise. Specifically, for each molecule, $100$ geometries are randomly divided into $80$ for training, $10$ for validation, and $10$ for testing. Here, the molecules in the test set are visible during training but the geometric structures are different from training structures.
QH9-dynamic-mol. In the QH9-dynamic-mol partitioning, QH9-dynamic-100k-mol and QH9-dynamic-300k-mol datasets utilize $999$ and $2,998$ molecules, respectively, and molecules are divided into training, validation, and testing subsets in a ratio of $0.8/0.1/0.1$. Importantly, different from the above QH9-dynamic-geo setup, all $100$ geometries corresponding to a specific molecule are grouped together and assigned to the same subset. This setup introduces a more challenging task than QH9-dynamic-geo since the geometries in the testing set correspond to different molecules as those in training.
@@ 4 Experiments
QH9-dynamic-100k-geo QHNet $149.62$ $92.88$ $96.85$ $834.47$ $94.45$
QH9-dynamic-100k-mol QHNet $416.99$ $153.68$ $173.92$ $9719.58$ $79.15$
QH9-dynamic-300k-geo a QHNet $166.99$ $95.25$ $100.19$ $843.14$ $94.95$
QH9-dynamic-300k-mol a QHNet $108.70$ $261.63$ $119.66$ $2178.15$ $90.72$
  • a
    The cost of training on QH9-dynamic-300k is similar compared to QH9-dynamic-100k, while it contains more data and achieves higher performance in molecule-wise split. Therefore, it is recommended to use QH9-dynamic-300k.
@@
Geometry-wise vs. molecule-wise generalization. We further explore geometry-wise and molecule-wise generalizability by analyzing the difficulty differences between the QH9-dynamic-geo and QH9-dynamic-mol tasks. We consider the results in Table 2 for these two tasks to be comparable given that both models are trained with a similar number of geometry structures. We note that the model in the QH9-dynamic-geo task demonstrates numerically better test performance than the model in the QH9-dynamic-mol task. This is consistent with our intention when designing the tasks. Specifically, in QH9-dynamic-geo, although the geometric structures in the test set are different, the molecules themselves are not entirely novel to the model due to the exposure during the training phase. In comparison, the QH9-dynamic-mol task presents a more challenging and demanding scenario. In particular, the test set geometries in QH9-dynamic-mol correspond to entirely different molecules than those seen during training. This task requires the model to generalize from its learned patterns to the unseen molecular structures. Notably, due to the increased number of trajectories with more diverse molecules in the QH9-dynamic-300k-mol dataset, the test performance significantly improves compared to the QH9-dynamic-100k-mol under the same training settings and cost. To summarize, both tasks serve as valuable testbeds in evaluating the model’s generalization ability, and our analysis shows that QH9-dynamic-mol task, which requires extrapolating to entirely new molecular structures, is notably more challenging and demanding.
Table 4: The performance of DFT calculation acceleration. Both models, trained on the QH9-stable-id split and the QH9-stable-ood split respectively, are evaluated on a common set of $50$ randomly chosen molecules from the intersection of their test sets, making their results directly comparable. Similarly, results from models trained on the QH9-dynamic-geo and QH9-dynamic-mol splits can be compared directly. This table demonstrates the possibility of using quantum tensor network as the initialization for quantum chemistry calculations, thereby speeding up the convergence. However, it is not necessary to run this metric for new designed models due to the computational cost of DFT.
@@
QH9-dynamic-100k-geo 1e Optimal ratio $0.056\scriptstyle{{{\color[rgb]{0.3,0.3,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,0.3}\pgfsys@color@rgb@stroke{0.3}{0.3}{0.3}\pgfsys@color@rgb@fill{0.3}{0.3}{0.3}\pm 0.006}}}$ QH9-dynamic-100k-mol 1e Optimal ratio $0.056\scriptstyle{{{\color[rgb]{0.3,0.3,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,0.3}\pgfsys@color@rgb@stroke{0.3}{0.3}{0.3}\pgfsys@color@rgb@fill{0.3}{0.3}{0.3}\pm 0.006}}}$
[…]
QH9-dynamic-300k-geo 1e Optimal ratio $0.057\scriptstyle{{{\color[rgb]{0.3,0.3,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,0.3}\pgfsys@color@rgb@stroke{0.3}{0.3}{0.3}\pgfsys@color@rgb@fill{0.3}{0.3}{0.3}\pm 0.007}}}$ QH9-dynamic-300k-mol 1e Optimal ratio $0.057\scriptstyle{{{\color[rgb]{0.3,0.3,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,0.3}\pgfsys@color@rgb@stroke{0.3}{0.3}{0.3}\pgfsys@color@rgb@fill{0.3}{0.3}{0.3}\pm 0.007}}}$
@@ Appendix A Appendix
Statistic Type QH9-stable-id QH9-stable-ood QH9-dynamic-300k-geo QH9-dynamic-300k-mol
# of atoms Mean 18.02 / 18.02 / 18.03 16.97 / 21.25 / 23.67 18.04 / 18.04 / 18.04 18.01 / 18.15 / 18.11
Median 18 / 18 / 18 17 / 21 / 23 18 / 18 / 18 18 / 18 / 18
Max 29 / 29 / 29 20 / 22 / 29 27 / 27 / 27 27 / 25 / 25
Min 3 / 6 / 4 3 / 21 / 23 7 / 7 / 7 7 / 10 / 9
# of electronics Mean 65.89 / 65.90 / 65.86 64.98 / 68.83 / 70.57 65.88 / 65.88 / 65.88 65.91 / 65.71 / 65.78
@@
Max 74 / 74 / 74 74 / 74 / 74 74 / 74 / 74 74 / 74 / 72
Min 10 / 18 / 24 10 / 56 / 58 24 / 24 /24 24 / 34 / 50
Size of Hamiltonian Mean 141.59 / 141.62 / 141.56 138.99 / 149.85 / 155.09 141.59 / 141.59 / 141.59 141.58 / 141.18 / 142.03
Median 144 / 144 / 144 142 / 150 / 154 144 / 144 / 144 144 / 144 / 144
Max 166 / 166 / 166 148 / 152 / 166 162 / 162 / 162 162 / 158 / 158
Min 18 / 36 / 48 18 / 126 / 130 54 / 54 / 54 54 / 72 / 102
Statistic Type QH9-dynamic-100k-geo QH9-dynamic-100k-mol
# of atoms Mean 16.53 / 16.53 / 16.53 16.56 / 16.38 / 16.38
Median 17 / 17 / 17 17 / 17 / 17
Max 19 / 19 / 19 19 / 19 / 19
Min 10 / 10 / 10 10 / 10 / 10
# of electronics Mean 64.68 / 64.68 / 64.68 64.71 / 64.44 / 64.68
Median 66 / 66 / 66 66 / 66 / 66
Max 74 / 74 / 74 74 / 70 / 70
Min 38 / 38 / 38 40 / 38 / 46
Size of Hamiltonian Mean 137.96 / 137.96 / 137.96 138.04 / 137.34 / 137.88
Median 140 / 140 / 140 142 / 142 / 142
Max 146 / 146 /146 146 / 146 / 146
Min 82 / 82 / 82 86 / 82 / 92
@@ A.1 Detailed dataset statistics
A.2 The analysis of the MD trajectory
To analyze the MD trajectory, we provide the MSD between the geometry in the initial structure and the geometry at each time step in Figure 1(a), and find that molecular geometry is changing over time. Furthermore, in Figure 1(b), we showcase the MAE of the Hamiltonian for initial structure and Hamiltonian compared to that for other geometries along the same trajectory. Note that the MAE error of the different geometries is much larger than the prediction error.
(a) The mean square distance (MSD) between the initial geometry and the geometry in the MD trajectory. Here it demonstrates the first MD trajectory in QH9-dynamic-300k.
(b) The mean absolute error (MAE) between the Hamiltonian matrix for the initial geometry and the Hamiltonian matrix for the geometry in the MD trajectory in QH9-dynamic-300k.
A.3 Test performance over molecule size
@@
A.4 Test performance over various training sets